what is the solution of the system of linear equations.... y=1/2x+8 and 2x+y=18
Answer by algebrahouse.com(1659)

"what is the solution of the system of linear equations.... y=1/2x+8 and 2x+y=18"

y = 1/2x + 8
2x + y = 18

2x + y = 18 {second equation}
2x + 1/2x + 8 = 18 {substituted 1/2x + 8, in for y, into 2nd equation}
5/2x + 8 = 18 {combined like terms}
5/2x = 10 {subtracted 8 from both sides}
x = 4 {divided both sides by 5/2...multiplied by reciprocal, 2/5}

y = 1/2x + 8 {top equation}
y = 1/2(4) + 8 {substituted 4, in for x, into y = 1/2x + 8}
y = 2 + 8 {multiplied 1/2 by 4}
y = 10 {combined like terms}

x = 4 and y = 10
www.algebrahouse.com